OpenStudy (anonymous):

How many liters of a 30% acid solution should be mixed with an 80% acid solution to get 70 L of a solution that is 50% acid?

OpenStudy (jdoe0001):

\(\large {\begin{array}{rccccclll} &liters&acidity&total\ liters \\\hline\\ 30\% \ sol.&a&0.30&0.30\cdot a\\ 60\%\ sol.&b&0.80&0.80\cdot b\\ \\\hline\\ mixture& a+b=70&0.50&70\cdot 0.50=35 \end{array} \\ \quad \\ \qquad thus\qquad a+b=70\implies a=70-b\qquad then \\ \quad \\ \begin{array}{rccccclll} &liters&acidity&total\ liters \\\hline\\ 30\% \ sol.&{\color{brown}{ 70-b}}&0.30&0.30\cdot ({\color{brown}{ 70-b}})\\ 60\%\ sol.&b&0.80&0.80\cdot b\\ \\\hline\\ mixture& a+b=70&0.50&70\cdot 0.50=35 \end{array} \\ \quad \\ then\qquad 0.30( {\color{brown}{ 70-b}})+0.08b=35 }\) solve for "b" to see how much you'd need of the 80% acid and to get "a", well, a = 70-b

OpenStudy (jdoe0001):

hmm kinda put 60%... should be 80% anyhow

OpenStudy (jdoe0001):

\(\large {\begin{array}{rccccclll} &liters&acidity&total\ liters \\\hline\\ 30\% \ sol.&{\color{brown}{ 70-b}}&0.30&0.30\cdot ({\color{brown}{ 70-b}})\\ 80\%\ sol.&b&0.80&0.80\cdot b\\ \hline\\ mixture& a+b=70&0.50&70\cdot 0.50=35 \end{array} \\ \quad \\ then\qquad 0.30( {\color{brown}{ 70-b}})+0.08b=35 }\)

OpenStudy (anonymous):

Okay, I did .30(70-y) + .80y=35, but am still not coming up with the same answer. They show 28, 12, 58, and 42 as one of the answers and I can't seem to find any of them.

OpenStudy (jdoe0001):

well, review your simplification, I can see the choices matching up

OpenStudy (anonymous):

I see. After dividing I came up with 28. Would I then say 70-28 to get my final answer?

OpenStudy (jdoe0001):

yeap